# Is GPT-OSS Good? A Comprehensive Evaluation of OpenAI’s Latest Open Source Models thanks:

### v1

@@ Abstract
In August 2025, OpenAI released GPT-OSS models, its first open weight large language models since GPT-2 in 2019, comprising two mixture of experts architectures with 120B and 20B parameters. We evaluated both variants against six contemporary open source large language models ranging from 14.7B to 235B parameters, representing both dense and sparse designs, across ten benchmarks covering general knowledge, mathematical reasoning, code generation, multilingual understanding, and conversational ability. All models were tested in unquantised form under standardised inference settings, with statistical validation using McNemar’s test and effect size analysis. Results show that gpt-oss-20B consistently outperforms gpt-oss-120B on several benchmarks, such as HumanEval and MMLU, despite requiring substantially less memory and energy per response. Both models demonstrate mid-tier overall performance within the current open source landscape, with relative strength in code generation and notable weaknesses in multilingual tasks. These findings provide empirical evidence that scaling in sparse architectures may not yield proportional performance gains, underscoring the need for further investigation into optimisation strategies and informing more efficient model selection for future open source deployments.
Index Terms:
Large language models, gpt-oss, model evaluation, benchmarking, reasoning models, mixture of experts, performance analysis
@@ I Introduction
The past five years have seen a rapid expansion of open source large language models (LLMs), with dense architectures such as Llama and Gemma joined by increasingly capable mixture of experts (MoE) designs. MoE architectures activate only a fraction of parameters per token [1, 2], offering a theoretical path to improved compute efficiency and scalability [3, 4, 5]. In August 2025, OpenAI released GPT OSS models, its first open weight models since GPT-2 in 2019 [6], introducing MoE variants with 20B and 120B total parameters into a landscape already shaped by competitive offerings from Meta [7], Google [8], DeepSeek [9], Alibaba [10, 11] and Microsoft [12].
@@
While scaling laws have been widely used to guide LLM development, empirical studies on their validity for sparse architectures remain limited. Existing evaluations have often focused on single domains or isolated benchmarks [13, 14], leaving open the question of whether larger MoE variants consistently outperform smaller ones across diverse capabilities. This gap matters both for understanding the training dynamics of sparsely activated networks and for making informed deployment choices under resource constraints.
@@
Our analysis serves two purposes. From a research perspective, it tests whether observed performance trends align with established scaling expectations in the MoE setting. From a practical perspective, it examines the cost performance profile of GPT-OSS relative to strong open source baselines, incorporating memory, latency, and energy measurements that are often omitted from academic reports. The results, presented in the following sections, provide evidence relevant to both theoretical modelling of scaling behaviour and the design of efficient deployment strategies in the evolving open source community.
@@ II Related Work
II-A Large Language Models Evaluation
The evaluation of large language models has evolved significantly since the introduction of transformer architectures. Early work by Radford et al. [6] laid the foundation for systematic LLM evaluation through the development of benchmarks like WebText and subsequent iterations. The field has since developed comprehensive evaluation frameworks addressing multiple dimensions of model capabilities.
Hendrycks et al. [15] introduced MMLU (Massive Multitask Language Understanding) as a comprehensive benchmark spanning 57 subjects, providing broad coverage of human knowledge domains. This benchmark has become a standard metric for assessing general understanding capabilities, though recent work highlights potential limitations including dataset contamination and the need for more challenging evaluation scenarios.
The evaluation of mathematical reasoning capabilities has been extensively studied through benchmarks like GSM8K, introduced by Cobbe et al. [16], which tests multi-step problem-solving abilities. Subsequent work by Lewkowycz et al. [17] with Minerva demonstrated that specialised training on mathematical content can significantly improve performance, raising questions about the generalisation of reasoning capabilities across domains.
II-B Code Generation Assessment
The HumanEval benchmark, developed by Chen et al. [18], established a rigorous framework for evaluating code generation capabilities through functional correctness testing. This work has been extended by Austin et al. [19] with MBPP (Mostly Basic Python Problems), and by Li et al. [20] with additional multilingual code generation benchmarks. Recent studies by Nijkamp et al. [21] demonstrate that code-specific pretraining significantly impacts performance, suggesting that general purpose models may face inherent limitations in programming tasks. The emergence of specialised coding models like Codex [18] and CodeGen [21] has prompted renewed interest in understanding the relationship between model architecture and code generation capabilities. Fried et al. [22] showed that instruction-tuned models can achieve competitive performance with significantly fewer parameters, challenging assumptions about scale requirements for code generation tasks.
II-C Multilingual Model Evaluation
Cross-lingual evaluation has gained prominence with the increasing deployment of LLMs in global contexts. The C-Eval benchmark, introduced by Huang et al. [23], provides comprehensive coverage of Chinese language understanding across multiple disciplines. Similar efforts include MMLU variants for other languages and the XTREME benchmark suite for cross-lingual transfer evaluation.
Research by Winata et al. [24] demonstrates that multilingual capabilities often come at the cost of monolingual performance, a phenomenon termed the “curse of multilinguality.” This trade-off is particularly relevant for models like GPT-OSS that aim for broad language coverage without language-specific optimisation.
II-D Mixture of Experts Architectures
The mixture of experts paradigm has experienced renewed interest with models like Switch Transformer by Fedus et al. [2] and GLaM by Du et al. [25], demonstrating that sparse models can achieve competitive performance with reduced computational requirements. The routing mechanisms and expert specialisation patterns in these models have been extensively studied by Lewis et al. [26], revealing complex dynamics between model capacity and task performance.
Recent work, such as the branch train merge approach from Artetxe et al. [27] and the Expert Choice routing from Zhou et al. [28], has advanced our understanding of optimal MoE configurations. These studies provide context for understanding the performance characteristics of GPT-OSS’s MoE implementation, particularly the unexpected inverse scaling observed between the 120B and 20B variants.
II-E Benchmark Limitations and Criticisms
Critical examination of evaluation methodologies has revealed significant limitations in current benchmarking practices. Bowman [29] argues that static benchmarks suffer from dataset contamination and overfitting, proposing dynamic evaluation protocols that adapt over time. The work of Raji et al. [30] highlights biases in benchmark construction that may favour certain model architectures or training approaches. The challenge of evaluating emergent capabilities not captured by traditional benchmarks has been addressed by Wei et al. [13], who demonstrate that many important model behaviours only manifest at certain scales or with specific prompting strategies. This work suggests that our evaluation of GPT-OSS models may not fully capture their potential, particularly in domains requiring complex reasoning or creativity.
II-F Open Source Model Ecosystem
The open source LLMs ecosystem has expanded dramatically following the release of models like LLaMA by Touvron et al. [7, 31] and Falcon by Almazrouei et al. [32]. These releases have enabled extensive community-driven research and development, leading to rapid improvements in model capabilities and efficiency. The work of Dettmers et al. [33] on quantisation techniques has made large models accessible to researchers with limited computational resources.
The impact of open source models on research velocity has been documented by Bommasani et al. [34], who argue that accessible models accelerate innovation while raising concerns about misuse and safety. The release of GPT-OSS represents a significant addition to this ecosystem, potentially shaping future development and research priorities.
II-G Performance Scaling Laws
Kaplan et al. [3] established foundational scaling laws relating model size, dataset size, and compute budget to model performance. Subsequent work by Hoffmann et al. [5] with Chinchilla challenged these assumptions, showing that many large models had been significantly undertrained. These findings have important implications for interpreting the performance of GPT-OSS models, particularly the unexpected inverse scaling between the 120B and 20B variants. Research by Schaeffer et al. [14] on emergent abilities suggests that performance scaling is more complex than initially understood. The presence of phase transitions and non-monotonic scaling in certain capabilities complicates predictions of model behaviour across scales.
II-H Advances in Evaluation Methodology
Work by Liang et al. [35] has advanced the development of rigorous evaluation methodologies through the HELM framework, which provides standardised protocols for comprehensive model assessment. The importance of statistical significance testing and effect size reporting has been emphasised by Card et al. [36], addressing reproducibility concerns in NLP research.
Recent work on contamination detection by Sainz et al. [37] highlights the need for careful experimental design and transparent reporting. Our methodology incorporates these considerations through careful sampling, statistical analysis, and comprehensive documentation of experimental procedures.
II-I Conversational Evaluation
The evaluation of conversational capabilities has evolved from single-turn assessment to complex multi-turn dialogue evaluation. The MT-Bench framework by Zheng et al. [38] provides a structured evaluation of conversational coherence and helpfulness, while work by Thoppilan et al. [39] on LaMDA demonstrates the importance of safety and factual grounding in dialogue systems.
Research by Roller et al. [40] on dialogue model evaluation highlights the challenge of balancing objectives, including engagement, consistency and informativeness. These considerations are particularly relevant for assessing GPT-OSS models’ conversational capabilities, where we observe trade-offs between response quality and computational efficiency.
@@ III-A Model Selection
Our evaluation encompasses eight contemporary large language models, selected to represent diverse architectural approaches, parameter scales, and development philosophies. The selection criteria, informed by recent survey work [41, 42], prioritized: (1) availability through standardised interfaces [43], (2) representation of major research institutions and companies, (3) variation in model sizes from 14.7B to 235B parameters, and (4) inclusion of both dense and sparse architectures [2, 44]. Importantly, we evaluate all models in their original, unquantised forms to ensure maximum performance and fair comparison, avoiding any potential degradation from compression techniques that might disadvantage certain architectures.
The evaluated models span multiple architectural paradigms. GPT-OSS models employ MoE architectures similar to Switch Transformer [2] and GShard [44], while dense models like Llama 3.3 follow the traditional transformer architecture [45]. Recent work by Rae et al. [46] and Chowdhery et al. [47] has demonstrated that both approaches can achieve strong performance, though with different computational trade-offs. Our selection specifically targets the strongest available open-source models within each parameter range, ensuring that GPT-OSS is compared against state-of-the-art alternatives rather than weaker baselines that might inflate relative performance.
@@
III-B Benchmark Selection
We employ ten complementary benchmark suites spanning diverse cognitive capabilities. The Massive Multitask Language Understanding (MMLU) benchmark [15] evaluates broad knowledge across 57 subjects, while GSM8K [16] assesses mathematical reasoning following methodologies established by Lewkowycz et al. [17]. Code generation capabilities are measured using HumanEval [18], with evaluation protocols refined by recent work from Li et al. [20]. Domain-specific evaluation includes FinQA[48] for financial reasoning, MedQA[49] for medical knowledge assessment, LegalQA[50] for legal understanding, and SciQ[51] for scientific reasoning. Common-sense reasoning is evaluated through PIQA (Physical Interaction Question Answering)[52], while DialogSum[53] assesses conversational summarisation capabilities. Cross-lingual evaluation employs C-Eval [23] for Chinese comprehension, complemented by methodologies from XTREME [54] and mT5 [55]. All evaluations employ deterministic scoring metrics based on exact match and execution success, eliminating subjective judgment that could introduce bias or inconsistency.
III-C Evaluation Protocol
Our evaluation protocol builds upon established best practices from the HELM framework [35] and BIG-bench [56]. Sampling strategies follow power analysis guidelines from Card et al. [36] to ensure statistical validity, while prompt engineering incorporates insights from Brown et al. [57], Wei et al. [58], and Kojima et al. [59]. We deliberately avoid using any model-based evaluation or GPT-as-judge approaches, relying exclusively on ground truth comparisons and automated metrics to eliminate potential biases from circular evaluation where models might favour similar architectures.
Figure 2: Parameter-performance relationship. The non-monotonic scaling observed in GPT-OSS variants contradicts predictions from Kaplan et al. [3] and suggests architectural inefficiencies.
Response processing employs techniques from Holtzman et al. [60] for decoding and Welleck et al. [61] for consistency checking. Temperature scaling follows recommendations from Ruis et al. [62], with nucleus sampling parameters based on findings from Meister et al. [63].
@@ III-D Statistical Analysis
Statistical rigour follows guidelines established by Dror et al. [64] and Benavoli et al. [65] for NLP evaluation. We employ McNemar’s test for paired comparisons [66] with Bonferroni correction [67], effect size calculation using Cohen’s d [68], and bootstrap confidence intervals following Efron and Tibshirani [69].
Recent work by Pimentel et al. [70] and Deutsch et al. [71] has highlighted the importance of proper statistical testing in NLP, particularly when comparing models with similar performance levels. Our analysis incorporates these recommendations, including variance estimation techniques from Koehn [72] and multiple testing corrections from Benjamini and Hochberg [73].
III-E Quality Assurance
Reproducibility measures follow recommendations from Pineau et al. [74] and Dodge et al. [75], including fixed random seeds, version locking, and comprehensive logging. Data contamination detection employs methods from Sainz et al. [37] and Magar and Schwartz [76], while evaluation gaming prevention follows guidelines from Dekoninck et al. [77]. To address concerns about benchmark overfitting, we include both established benchmarks and recent evaluation sets released after the training cutoff of evaluated models, ensuring that performance differences reflect genuine capabilities rather than memorisation.
III-F Limitations and Ethical Considerations
Our methodology acknowledges limitations identified in prior benchmark critiques [29, 30, 78]. Ethical considerations follow guidelines from Bender et al. [79] and Birhane et al. [80], including assessment of potential biases and environmental impact following Strubell et al. [81] and Patterson et al. [82]. While critics might argue that benchmarks inadequately capture real-world performance, our comprehensive suite spanning knowledge, reasoning, code, and multilingual tasks provides robust evidence of relative model capabilities. The consistency of performance patterns across diverse evaluation domains strengthens confidence in our findings beyond what any single benchmark could provide.

### v3

@@ Abstract
In August 2025, OpenAI released GPT-OSS models, its first open weight large language models since GPT-2 in 2019, comprising two mixture of experts architectures with 120B and 20B parameters. We evaluated both variants against six contemporary open source large language models ranging from 14.7B to 235B parameters, representing both dense and sparse designs, across ten benchmarks covering general knowledge, mathematical reasoning, code generation, multilingual understanding, and conversational ability. All models were tested in unquantised form under standardised inference settings, with statistical validation using McNemar’s test and effect size analysis. Results show that gpt-oss-20B consistently outperforms gpt-oss-120B on several benchmarks, such as HumanEval and MMLU, despite requiring substantially less memory and energy per response. Both models demonstrate mid-tier overall performance within the current open source landscape, with relative strength in code generation and notable weaknesses in multilingual tasks. These findings provide empirical evidence that scaling in sparse architectures may not yield proportional performance gains, underscoring the need for further investigation into optimisation strategies and informing more efficient model selection for future open source deployments. More details and evaluation scripts are available at the Project Webpage.
@@ I Introduction
The past five years have seen a rapid expansion of open source large language models (LLMs), with dense architectures such as Llama and Gemma joined by increasingly capable mixture of experts (MoE) designs. MoE architectures activate only a fraction of parameters per token [shazeer2017outrageously, fedus2022switch], offering a theoretical path to improved compute efficiency and scalability [kaplan2020scaling, henighan2020scaling, hoffmann2022training]. In August 2025, OpenAI released GPT OSS models, its first open weight models since GPT-2 in 2019 [radford2019language], introducing MoE variants with 20B and 120B total parameters into a landscape already shaped by competitive offerings from Meta [touvron2023llama], Google [gemma2024], DeepSeek [deepseek2024], Alibaba [bai2023qwen, qwen2024technical] and Microsoft [abdin2024phi].
@@
While scaling laws have been widely used to guide LLM development, empirical studies on their validity for sparse architectures remain limited. Existing evaluations have often focused on single domains or isolated benchmarks [wei2022emergent, schaeffer2024emergent], leaving open the question of whether larger MoE variants consistently outperform smaller ones across diverse capabilities. This gap matters both for understanding the training dynamics of sparsely activated networks and for making informed deployment choices under resource constraints.
@@
Our analysis serves two purposes. From a research perspective, it tests whether observed performance trends align with established scaling expectations in the MoE setting. From a practical perspective, it examines the cost performance profile of GPT-OSS relative to strong open source baselines, incorporating memory, latency, and energy measurements that are often omitted from academic reports. As illustrated in Fig 1 and elaborated in subsequent sections, these results provide evidence relevant to both theoretical modelling of scaling behaviour and the design of efficient deployment strategies in the evolving open source community.
@@ II Related Work
II-A Evaluation of Large Language Models
The evaluation of large language models (LLMs) has broadened as models have grown in scale and capability. Early work by Radford et al. [radford2019language] laid the foundation for systematic assessment with initial benchmarks such as WebText [radford2019language]. Hendrycks et al. [hendrycks2021measuring] introduced MMLU, covering 57 subjects, which became a standard for measuring broad knowledge. For mathematical reasoning, Cobbe et al. [cobbe2021training] proposed GSM8K to evaluate multi-step arithmetic problems, and Lewkowycz et al. [lewkowycz2022solving] showed with Minerva [lewkowycz2022solving] that domain-specific training can dramatically improve mathematical performance and raising questions about cross domain generalisation.
Code generation has emerged as another central evaluation domain. Chen et al. [chen2021evaluating] introduced HumanEval to test functional correctness in programming tasks. Austin et al. [austin2021program] and Li et al. [li2023starcoder] extended this with MBPP and multilingual coding benchmarks. Studies have demonstrated that pretraining on large code corpora yields substantial gains [nijkamp2022codegen], and Fried et al. [fried2022incoder] showed that instruction-tuned smaller models can rival much larger systems, challenging assumptions about scale requirements in coding tasks.
Beyond English, evaluation efforts have expanded to multilingual contexts. Huang et al. [huang2023c] introduced C-Eval, a large-scale benchmark for Chinese, while XTREME and MMLU variants extended coverage to other languages. These studies revealed the so-called “curse of multilinguality” [winata2021language], where expanding language coverage often reduces monolingual performance. This trade-off is directly relevant to GPT-OSS, which aims to support many languages without language-specific optimisation.
Conversational ability is now assessed with benchmarks designed for multi-turn dialogue. Zheng et al. [zheng2023judging] proposed MT-Bench to measure coherence and helpfulness, while Thoppilan et al. [thoppilan2022lamda] stressed the importance of factual grounding and safety in open domain dialogue. Roller et al. [roller2021recipes] highlighted the difficulty of balancing engagement, consistency and informativeness. These considerations frame our assessment of GPT-OSS, where conversational quality is weighed against computational efficiency.
II-B Model Architectures and Scaling
Architecture choices strongly influence model performance and efficiency. Mixture of Experts (MoE) architectures, which activate only a subset of parameters per token, have become a focus due to their efficiency advantages. The Switch Transformer [fedus2022switch] and GLaM [du2022glam] showed that sparse models can rival dense models with significantly reduced compute. Work on routing and expert allocation [lewis2020retrieval] revealed complex trade-offs, and methods such as Branch-Train-Merge [artetxe2021efficient] and Expert-Choice routing [zhou2022mixture] improved training stability and performance. GPT-OSS adopts an MoE design, and our findings reveal an unusual outcome: the 20B model sometimes surpasses the 120B model, highlighting an inverse scaling effect observed in other MoE systems.
Scaling laws offer another perspective. Kaplan et al. [kaplan2020scaling] proposed power law relationships between model size, data and compute, suggesting predictable improvements with scale. Hoffmann et al. [hoffmann2022training] refined this view, showing many large models had been under-trained and introducing the Chinchilla scaling approach. More recently, Schaeffer et al. [schaeffer2024emergent] demonstrated that performance scaling can be non-monotonic, with phase transitions and regressions at certain sizes. These findings complicate the interpretation of GPT-OSS, where the smaller model sometimes outperforms the larger, underscoring that parameter count alone does not guarantee progress.
II-C Open Source Community and Evaluation
Open source model releases have reshaped the field. Llama [touvron2023llama, touvron2023llama2] and Falcon [almazrouei2023falcon] gave the community access to strong baselines, and techniques such as post-training quantisation [dettmers2022gpt3] lowered deployment costs, accelerating iteration [bommasani2021opportunities]. Since 2024, both proprietary and open initiatives have broadened the landscape with Gemini 2.5 [comanici2025gemini], Claude Opus 4.1 [anthropic2025claude4], Qwen3 [yang2025qwen3] and the latest DeepSeek-V3.1 [deepseek2024, deepseek2025v3.1]. Within this setting, GPT-OSS contributes MoE-based models at 20B and 120B scales, offering a controlled case study of sparse scaling in comparison with contemporary selected open source models.
Benchmarking practices, however, face persistent challenges. Static leaderboards encourage overfitting and contamination [bowman2023eight], while benchmark construction may embed biases [raji2021ai]. Certain abilities emerge only at larger scales or with specific prompting [wei2022emergent], limiting the coverage of fixed tests. To address these issues, Liang et al. [liang2022holistic] proposed the HELM framework for holistic evaluation, and Card et al. [card2020little] emphasised statistical validity and effect size reporting. Efforts to detect dataset overlap [sainz2023nlp] further highlight the need for transparency. Our evaluation of GPT-OSS follows these principles by combining diverse benchmarks, applying significance testing and reporting effect sizes, and documenting procedures in detail, ensuring fair comparison and reproducibility.
@@ III Methodology
Distinguished from prior evaluation frameworks, our study directly compares mixture-of-experts (MoE) and dense models under the same experimental baseline. All models are tested in their original full-precision form to avoid biases from quantization or compression. This setup ensures fair and reproducible comparison across architectures, parameter scales, and benchmark tasks.
@@ III-A Model Selection
Our evaluation encompasses eight contemporary large language models, selected to represent diverse architectural approaches, parameter scales, and development philosophies. The selection criteria, informed by recent survey work [zhao2023survey, minaee2024large], prioritized: (1) availability through standardised interfaces [gerganov2023llama], (2) representation of major research institutions and companies, (3) variation in model sizes from 14.7B to 235B parameters, and (4) inclusion of both dense and sparse architectures [fedus2022switch, lepikhin2021gshard]. Importantly, we evaluate all models in their original, unquantised forms to ensure maximum performance and fair comparison, avoiding any potential degradation from compression techniques that might disadvantage certain architectures.
The evaluated models span multiple architectural paradigms. GPT-OSS models employ MoE architectures similar to Switch Transformer [fedus2022switch] and GShard [lepikhin2021gshard], while dense models like Llama 3.3 follow the traditional transformer architecture [vaswani2017attention]. Recent work by Rae et al. [rae2021scaling] and Chowdhery et al. [chowdhery2022palm] has demonstrated that both approaches can achieve strong performance, though with different computational trade-offs. Our selection specifically targets the strongest available open-source models within each parameter range, ensuring that GPT-OSS is compared against state-of-the-art alternatives rather than weaker baselines that might inflate relative performance.
III-B Benchmark Selection
We employ ten complementary benchmark suites spanning diverse cognitive capabilities. The Massive Multitask Language Understanding (MMLU) benchmark [hendrycks2021measuring] evaluates broad knowledge across 57 subjects, while GSM8K [cobbe2021training] assesses mathematical reasoning following methodologies established by Lewkowycz et al. [lewkowycz2022solving]. Code generation capabilities are measured using HumanEval [chen2021evaluating], with evaluation protocols refined by recent work from Li et al. [li2023starcoder]. Domain-specific evaluation includes FinQA[chen2022finqadatasetnumericalreasoning] for financial reasoning, MedQA[zhang2018medicalexamquestionanswering] for medical knowledge assessment, LegalQA[li2024legalqa] for legal understanding, and SciQ[welbl2017crowdsourcingmultiplechoicescience] for scientific reasoning. Common-sense reasoning is evaluated through PIQA (Physical Interaction Question Answering)[bisk2019piqareasoningphysicalcommonsense], while DialogSum[chen2021dialogsumreallifescenariodialogue] assesses conversational summarisation capabilities. Cross-lingual evaluation employs C-Eval [huang2023c] for Chinese comprehension, complemented by methodologies from XTREME [hu2020xtreme] and mT5 [xue2021mt5]. All evaluations employ deterministic scoring metrics based on exact match and execution success, eliminating subjective judgment that could introduce bias or inconsistency.
III-C Evaluation Protocol
Our evaluation protocol builds upon established best practices from the HELM framework [liang2022holistic] and BIG-bench [srivastava2022beyond]. Sampling strategies follow power analysis guidelines from Card et al. [card2020little] to ensure statistical validity, while prompt engineering incorporates insights from Brown et al. [brown2020language], Wei et al. [wei2022chain], and Kojima et al. [kojima2022large]. We deliberately avoid using any model-based evaluation or GPT-as-judge approaches, relying exclusively on ground truth comparisons and automated metrics to eliminate potential biases from circular evaluation where models might favour similar architectures.
We implemented automated consistency checks to validate the alignment between questions and their corresponding ground truth labels across all datasets. This included verification of multiple-choice answer indexing, numerical precision for mathematical problems in GSM8K and FinQA, and code execution correctness for HumanEval examples. Any misalignments detected during this process were manually reviewed and corrected where possible.
@@ III-C Evaluation Protocol
Figure 2: Parameter-performance relationship. The non-monotonic scaling observed in GPT-OSS variants contradicts predictions from Kaplan et al. [kaplan2020scaling] and suggests architectural inefficiencies.
We use randomisation procedures to minimise systematic bias in our evaluation. Models are evaluated in random order within each benchmark to prevent ordering effects, and test examples are shuffled to eliminate position-dependent biases. This approach ensures that observed performance differences reflect genuine model capabilities rather than experimental artefacts.
To ensure the objectivity of our findings, particularly for benchmarks requiring manual quality checks, we implemented a blind evaluation protocol. This methodology is crucial for mitigating cognitive biases in human raters, such as confirmation bias or the expectancy effect, where prior knowledge of a model’s identity could influence assessment [wang2023large]. Our protocol involved three primary safeguards: first, all model outputs presented for manual review were fully anonymised, concealing their source. Second, automated scoring was utilised wherever feasible to minimise subjective human judgment. Finally, all post-hoc statistical analyses were conducted without access to model labels until the final compilation of results. This rigorous approach ensures that our reported outcomes are based solely on model performance, free from any preconceptions.
Response processing employs techniques from Holtzman et al. [holtzman2019curious] for decoding and Welleck et al. [welleck2019neural] for consistency checking. Temperature scaling follows recommendations from Ruis et al. [ruis2024goldilocks], with nucleus sampling parameters based on findings from Meister et al. [meister2023locally].
@@ III-D Statistical Analysis
Statistical rigour follows guidelines established by Dror et al. [dror2018hitchhiker] for NLP evaluation, through a systematic statistical approach with four key components.
Primary Statistical Test. We employ McNemar’s test for pairwise model comparisons [mcnemar1947note]. This test is specifically designed for comparing paired binary outcomes and accounts for the dependence between models evaluated on identical data.
Multiple Comparison Correction. Given the comparisons among 8 models, we apply Benjamini-Hochberg correction [benjamini1995controlling] to control the false discovery rate, providing greater statistical power than traditional Bonferroni correction while maintaining appropriate error control for multiple testing scenarios.
Effect Size Estimation. We report accuracy differences with their corresponding confidence intervals to quantify the practical magnitude of performance gaps. This approach provides interpretable measures of effect size appropriate for model comparison tasks [cohen1988statistical].
Uncertainty Estimation. Bootstrap confidence intervals under the guidelines of [efron1994introduction] provide robust uncertainty estimates for all performance metrics. We also follow the established practices in computational linguistics evaluation [koehn2004statistical].
Recent work emphasising proper statistical validation in NLP [pimentel2020information, deutsch2021statistical] also guided our approach, particularly for comparing models with similar performance levels where small differences may not reflect genuine capability gaps.